OpenStudy (anonymous):

What is the volume of a cone with a radius of 18 cm and a height of 10 cm? Use 3.14 for pi. Round your answer to the nearest tenth. 3275.4 cm3 3391.2 cm3 3540.8 cm3 4000.2 cm3

OpenStudy (anonymous):

do u know the formula?

OpenStudy (anonymous):

no

OpenStudy (briellshaw):

My dad says its the 3rd one

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ok so its \[V=\pi r^2 h/3 \]

OpenStudy (anonymous):

thank you

OpenStudy (briellshaw):

your welcome

OpenStudy (anonymous):

dont u want to understand it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

explain ??? pls

OpenStudy (anonymous):

ayeshaafzal221

OpenStudy (anonymous):

ok so the u use this formula

OpenStudy (anonymous):

\[V=\pi r^2 \frac{ h }{ 3 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

where r is radius and h is height

OpenStudy (anonymous):

\[V= \pi \times 18^2 \times \frac{ 10 }{ 3 }\]

OpenStudy (anonymous):

now if u put that in ur calculator, u ll get ur answer :)

OpenStudy (briellshaw):

I am sorry it was the 2nd one

OpenStudy (briellshaw):

this is what I do

OpenStudy (briellshaw):

18*18=324*3.14=1017.36*10/3=3391.2 So B. Would be the Correct answer again my bad for misguiding you we must have made a wrong turn in the equation

OpenStudy (anonymous):

thats why u should never just give out the answer , work through it first

OpenStudy (briellshaw):

If you read my reply I clearly said i did the math and took a wrong turn and i got c